# Evidence for large baryonic feedback at low and intermediate redshifts from kinematic Sunyaev-Zel’dovich observations with ACT and DESI photometric galaxies

### v1

@@ V Summary and implications
In this paper, we present the highest SNR measurement of the gas profiles around galaxy groups using the kSZ effect: we detect the signal at 13 and find that it deviates from the dark matter at 40 (see Fig. 1 and Table 1). Such strong baryonic feedback exceeds predictions from state-of-the-art hydrodynamical simulations such as IllustrisTNG (TNG300-1). It also suggests that baryons might play a more significant role than assumed in many cosmological analyses and alleviate tensions such as ‘Lensing is low’ and ‘Low .’ Properly accounting for baryonic feedback is critical to placing robust constraints on many open questions, including the mass of neutrinos and the nature of dark matter – questions that will be crucial to future cosmological surveys such as Roman, Euclid, and the Vera Rubin Observatory. Our measurements of the gas profiles can be used to calibrate the subgrid models of hydrodynamical simulations, a task typically complicated by the low detection sensitivity to gas on the outskirts of halos (i.e., the ‘missing baryon’ problem). Combining kSZ measurements with tSZ and X-ray measurements, which provide access to additional quantities such as the temperature and cooling rate around galaxies [53, 54], will enable us to fully solve the gas thermodynamics of groups and clusters and shed light on the role of feedback in galaxy evolution.
To put our findings into perspective in relation to the tension, we offer a back-of-the-envelope calculation. The value of the parameter as measured by cosmic shear surveys is about 10% lower than the Planck values [14], suggesting a 20% suppression to the matter power spectrum relative to a dark-matter-only universe across a range of scales . Let’s consider a point halfway between these values, at . Assuming that the gas and dark matter follow a Gaussian profile [33], corresponds to . As seen from Fig. 2, at about half of the gas is missing relative to the dark matter at a significance of 6, which would manifest itself in the matter power spectrum as a 16% suppression with an uncertainty of 3% (assuming that baryons make up 16% of the total matter). This is roughly the effect of baryons needed to explain the tension, as suggested in Fig. 6 in Ref. [45] and Fig. 6 in Ref. [55], and ball-park matches the 20% suppression on the matter power corresponding to the measured value in cosmic shear experiments. Additionally, we note that our 13 detection of the gas can be converted into a 2.5% modeling error on the matter power spectrum (assuming a single fitting parameter for the amplitude), which is well beyond the required accuracy for future experiments (e.g., the Vera Rubin Observatory). As such, weak lensing studies taking into account kSZ measurements to model the effect of baryons will be especially powerful for future cosmic shear analyses [56].
@@
To make the kSZ measurement, we use the ACT temperature map and the DESI photometric galaxy sample of LRGs, which we split into four redshift bins. We detect the signal with respect to the dark matter in each of them at 10 (see Table 1). This allows us to study for the first time the redshift evolution of the baryonic feedback through kSZ stacking and place tighter constraints on the allowed astrophysical feedback models. The fact that we see little evolution of the signal suggests that the population of red galaxies probed by DESI is fairly stable and that the AGN feedback is of similar magnitude across . A major advantage of the gas profiles obtained using the kSZ is that they are practically systematics-free, as the velocity-weighted stacking we perform cancels additive contaminants such as the tSZ and CIB (see Fig. 3 and Schaan et al. [12]). Importantly, unlike many other astrophysical probes, the signal is directly proportional to the amount of gas and independent of other properties such as temperature. To put our findings into perspective, we compare the observed gas density profiles with mock measurements extracted from the state-of-the-art hydrodynamical simulation TNG300-1 and its predecessor Illustris-1. In particular, we mimic the stacking and LRG selection process of the observational analysis and test various scenarios related to the simulation targeting choices (satellite fraction, number density, halo mass, velocity uncertainty) to quantify the allowed variations (see Fig. 2). We find that the baryonic feedback in the TNG300-1 simulation is not sufficiently strong to push enough of the gas out of the halo virial radius, whereas Illustris-1 accomplishes that more successfully. Future work examining the origin of this is instrumental to reconciling the theory and observations of the gas-dark matter link.
@@ Acknowledgements.
Data points for the figures are available at https://doi.org/10.5281/zenodo.12633573. We would like to thank Minh Nguyen, Vid Irsic, Alex Amon, Daniel Gruen, and Yulin Gong for providing useful comments during the preparation of this manuscript. B.H. thanks the Miller Institute for financially supporting her postdoctoral research. S.F. is supported by Lawrence Berkeley National Laboratory and the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. B.R. and E.S. received support from the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. The Flatiron Institute is supported by the Simons Foundation. EC acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 849169). MM acknowledges support from NSF grants AST-2307727 and AST-2153201 and NASA grant 21-ATP21-0145. GSF acknowledges support through the Isaac Newton Studentship and the Helen Stone Scholarship at the University of Cambridge. GSF and FQ furthermore acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 851274). CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. RH acknowledges funding from the NSERC Discovery Grant RGPIN-2018-05750 and the Connaught Fund. This research was enabled in part by support provided by the SciNet cluster and the Digital Research Alliance of Canada. This work was supported by a grant from the Simons Foundation (CCA 918271, PBL). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall -band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/. Support for ACT was through the U.S. National Science Foundation through awards AST-0408698, AST-0965625, and AST-1440226 for the ACT project, as well as awards PHY-0355328, PHY-0855887 and PHY-1214379. Funding was also provided by Princeton University, the University of Pennsylvania, and a Canada Foundation for Innovation (CFI) award to UBC. ACT operated in the Parque Astronómico Atacama in northern Chile under the auspices of the Agencia Nacional de Investigación y Desarrollo (ANID). The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. Detector research at NIST was supported by the NIST Innovations in Measurement Science program. Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto. We thank the Republic of Chile for hosting ACT in the northern Atacama, and the local indigenous Licanantay communities whom we follow in observing and learning from the night sky.Appendix A Comparison with CMASS measurement

### v2

@@ V Summary and implications
In this paper, we present the highest SNR measurement of the gas profiles around galaxy groups using the kSZ effect: we detect the signal at 13 and find strong evidence of deviation from dark matter profiles, as well as a preference for larger feedback compared to some commonly used state-of-the-art hydrodynamical simulations such as IllustrisTNG.
@@
It also suggests that baryons might play a more significant role than assumed in many cosmological analyses and alleviate tensions such as ‘Lensing is low’ and ‘Low .’ Properly accounting for baryonic feedback is critical to placing robust constraints on many open questions, including the mass of neutrinos and the nature of dark matter – questions that will be crucial to future cosmological surveys such as Roman, Euclid, and the Vera Rubin Observatory. Our measurements of the gas profiles can be used to calibrate the subgrid models of hydrodynamical simulations, a task typically complicated by the low detection sensitivity to gas on the outskirts of halos (i.e., the ‘missing baryon’ problem). Combining kSZ measurements with tSZ and X-ray measurements, which provide access to additional quantities such as the temperature and cooling rate around galaxies [54], will enable us to fully solve the gas thermodynamics of groups and clusters and shed light on the role of feedback in galaxy evolution.
@@
To make the kSZ measurement, we use the ACT temperature map and the DESI photometric galaxy sample of LRGs, which we split into four redshift bins. We detect the signal with respect to the dark matter in each of them at 10 (see Table 1). This allows us to study for the first time the redshift evolution of the baryonic feedback through kSZ stacking and place tighter constraints on the allowed astrophysical feedback models. The fact that we see little evolution of the signal suggests that the population of red galaxies probed by DESI is fairly stable and that the AGN feedback is of similar magnitude across . A major advantage of the gas profiles obtained using the kSZ is that they are practically systematics-free, as the velocity-weighted stacking we perform cancels additive contaminants such as the tSZ and CIB (see App. A on null tests as well as [55]). Importantly, unlike many other astrophysical probes, the signal is directly proportional to the amount of gas and independent of other properties such as temperature. To put our findings into perspective, we compare the observed gas density profiles with mock measurements extracted from the state-of-the-art hydrodynamical simulation TNG300-1 and its predecessor, Illustris-1. In particular, we mimic the stacking and LRG selection process of the observational analysis and test various scenarios related to the simulation targeting choices (satellite fraction, number density, halo mass, velocity uncertainty) to quantify the allowed variations (see Fig. 2). We find that the baryonic feedback in the TNG300-1 simulation is not sufficiently strong to push enough of the gas out of the halo virial radius, whereas Illustris-1 accomplishes that more successfully. Future work examining the origin of this is instrumental to reconciling the theory and observations of the gas-dark matter link.
@@ Acknowledgements.
Data points for the figures are available at https://doi.org/10.5281/zenodo.12633573. We would like to thank Minh Nguyen, Vid Irsic, Alex Amon, Daniel Gruen, Yulin Gong, as well as the anonymous referees for providing very useful comments. B.H. thanks the Miller Institute for financially supporting her postdoctoral research. S.F. is supported by Lawrence Berkeley National Laboratory and the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. B.R. and E.S. received support from the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231. The Flatiron Institute is supported by the Simons Foundation. EC acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 849169). MM acknowledges support from NSF grants AST-2307727 and AST-2153201 and NASA grant 21-ATP21-0145. GSF acknowledges support through the Isaac Newton Studentship and the Helen Stone Scholarship at the University of Cambridge. GSF and FQ furthermore acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 851274). CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. CS acknowledges support from the Agencia Nacional de Investigación y Desarrollo (ANID) through Basal project FB210003. RH acknowledges funding from the NSERC Discovery Grant RGPIN-2018-05750 and the Connaught Fund. This research was enabled in part by support provided by the SciNet cluster and the Digital Research Alliance of Canada. This work was supported by a grant from the Simons Foundation (CCA 918271, PBL). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science, Innovation and Universities of Spain (MICIU/AEI/10.13039/501100011033), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation. The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall -band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/. Support for ACT was through the U.S. National Science Foundation through awards AST-0408698, AST-0965625, and AST-1440226 for the ACT project, as well as awards PHY-0355328, PHY-0855887 and PHY-1214379. Funding was also provided by Princeton University, the University of Pennsylvania, and a Canada Foundation for Innovation (CFI) award to UBC. ACT operated in the Parque Astronómico Atacama in northern Chile under the auspices of the Agencia Nacional de Investigación y Desarrollo (ANID). The development of multichroic detectors and lenses was supported by NASA grants NNX13AE56G and NNX14AB58G. Detector research at NIST was supported by the NIST Innovations in Measurement Science program. Computing for ACT was performed using the Princeton Research Computing resources at Princeton University, the National Energy Research Scientific Computing Center (NERSC), and the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by the CFI under the auspices of Compute Canada, the Government of Ontario, the Ontario Research Fund–Research Excellence, and the University of Toronto. We thank the Republic of Chile for hosting ACT in the northern Atacama, and the local indigenous Licanantay communities whom we follow in observing and learning from the night sky.
Appendix A Null test
In Fig. 3, we validate that our measured signal is indeed due to the kSZ effect rather than correlated contaminants such as the cosmic infrared background (CIB) or tSZ. For this test, we shuffle randomly the sample of reconstructed velocities to obtain 1000 realizations and perform the stacking of Eq. 4 but using the shuffled velocities on each galaxy. Since the true kSZ effect is proportional to the line-of-sight velocity, deleting that information in the shuffling should lead to a null measurement, which is indeed what we find. In particular, the null for the four redshift bins is 5.9, 5.0, 3.3 and 5.7, and the -value is 0.75, 0.83, 0.95, 0.77, respectively. The tSZ and CIB effects do not have such dependence666Foreground emission such as from the CIB is also Doppler-boosted and acquires a term proportional to velocity. However, we find that this contribution is about two orders of magnitude smaller than the kSZ signal [56], and we’ll therefore neglect it in this analysis., and hence if our ACT map is correctly cleaned, any additive contamination should also vanish. We note that this argument also holds true for the unshuffled (original) kSZ measurement, which is robust to additive contributions.
Appendix B Covariance matrix
The covariance matrix of the data is estimated via bootstrapping. In particular, for a given sample (e.g., the first redshift bin of the Main LRG sample), we generate 10,000 sets of kSZ measurements, where each set consists of a random draw (with repetition) from the kSZ measurements (equal to the number of galaxies) in that sample. We then compute the mean profile of each set and compute the covariance matrix across all 10,000 sets. We opt for 10,000 sets, as the error on the measurement error scales as the square root of that, and is thus about 1%. This produces an unbiased estimate of the covariance, in the limit of independent noise realizations from galaxy to galaxy. The assumption is not correct for large apertures where the temperature map cutouts (submaps) overlap, and the inferred covariance is only accurate at the 10% level (see [12]). However, given our uncertainty, this level of accuracy is sufficient.
In Fig. 7, we illustrate the stacked 2D kSZ signal for the Main DR9 sample of DESI LRGs. We apply a Wiener filter () that effectively high-pass filters the CMB temperature map to isolate the small-scale signal that is the most affected by the kSZ effect. We can see by eye the extended gas envelope of the DESI LRGs, which spans several arcmin. In particular, using TNG300 we extract the average virial radius of the DESI LRGs at the mean redshift of the sample, 1.75 arcmin. As the CMB map is convolved with a 1.6 arcmin (FWHM) Gaussian beam, we can estimate the beam-convolved mean virial radius to be arcmin. This roughly corresponds to the mean halo virial radius at that redshift.
Appendix C Simulations
To put our findings into context, we compare the gas density anisotropy we observe in the data against two hydrodynamical simulations: IllustrisTNG [47] and Illustris [48], which have very different AGN and supernova feedback prescriptions. To make the measurements in the simulations, we construct 2D maps of kSZ, then convolve them with a 1.6 arcmin Gaussian (corresponding to the DR6 map beam) to roughly match the ACT beam effects, and finally perform the stacking in an equivalent fashion as done to the data.
Each simulation outputs a number of useful quantities: gas mass (), electron abundance (), and internal energy () for each gas particle . Assuming a primordial hydrogen mass fraction of , we compute the volume-weighted electron number density, as
where the proton mass and the volume of the gas cell. We then compute the 2D maps of the kSZ and the optical depth by binning the gas particles into a (10000, 10000) grid, so that the optical depth in cell is given by:
and the momentum of the electron density is:
where is the area of each grid cell (of size 0.01 Mpc), is the Thomson cross section, and is the speed of light. Analogously, we can compute the Y-Compton map relevant to the tSZ effect and the Thomson optical depth map relevant to e.g., patchy screening.
We select LRG-like galaxies in the two simulations via an abundance matching approach and stack them at their locations on 2D CMB data mimicking ACT maps.Namely, we rank-order the galaxies by stellar mass and select the top such that, for the Main sample and for the Extended sample. We find that the mean halo mass of LRGs in TNG300-1 matches very well the inferred mean halo mass of DESI LRGs at , , and at , [43]. We then perform 2D stacking on the optical depth maps at the positions of these galaxy samples.
Appendix D Effect of photometric noise and recovery of the true
In this appendix, we explore two important effects and show that they are properly accounted for in our analysis. First, we study the effect of reconstruction on the shape of the CAP kSZ profiles. To this end, we repeat the CAP kSZ measurement detailed in Section III on an LRG-like galaxy sample, where we a) use the true host halo velocities in the stacking and b) use the reconstructed velocities with 2% uncertainties on the photometric redshifts (similar to DR9 and DR10 used in this analysis). As in the case of the DESI data, we perform reconstruction by adopting the package pyrecon. Our simulation mocks aim to mimic the DESI LRG selection and the ACT CMB map as closely as possible. In particular, we employ the full-sky AbacusSummit simulation, AbacusSummit_huge_c000_ph201, and populate its halo light cones with an HOD matching the properties of the Extended LRG sample (for more details, see Refs. [57, 33]). We then add redshift space distortions to each galaxy and photometric noise of 2%. We cut the galaxy catalog into the DESI footprint and perform reconstruction using the same settings in pyrecon as in the real data.
On the CMB side, we generate the optical depth field on the AbacusSummit density shells utilizing the ‘Transfer Function’ method of Ref. [58, 59]. When creating the kSZ maps, we multiply each optical depth shell by the line-of-sight velocity shell before integrating (i.e., summing up the shells) along the line-of-sight. Finally, we reproject the HEALPiX map into the shape and footprint of the ACT DR6 map using the package pixell. We can now treat these two data sets (the galaxy catalog and the kSZ/ map) as if they are ‘real data’ and perform the corresponding kSZ analysis. We correct the reconstructed profile by normalizing the estimator in Eq. 4 to account for the imperfect reconstruction, with , appropriate for the photometric redshift error adopted here [33]. To put these findings into the context of the present work, we show the error bars from our main result, Fig. 2.
As seen from Fig. 5, all three curves are in excellent agreement with each other and well within the measurement error bars. We find no evidence for scale dependence (0.1 difference between the two) of the coefficient despite the presence of photometric noise and redshift space distortions, verifying our methodology. This is expected because the cosmological velocity field has a correlation length much larger than a typical gas halo, and confirms the analytical results of [60]. Furthermore, since the first version of this paper, the corresponding analysis for the DESI Y1 spectroscopic sample has appeared [45]. A detailed comparison between the spectroscopic and photometric analyses is contained in that reference, but here we note that the excellent agreement between the two further confirms that photometric redshift errors do not bias the signal or introduce a scale dependence, beyond changing the value of as expected.
The second takeaway from this plot is that the reconstructed curve recovers the true optical depth curve very well, demonstrating that our kSZ measurements indeed recover the true optical depth, as intended.
Appendix E Mass evolution
Understanding the mass evolution of baryonic feedback is key to properly modeling its effect on cosmological observables such as cosmic shear, and the SZ effect is a versatile probe capable of capturing multiple gas properties [see e.g., 61, 16, 62, 63, 64, 65, for novel probes and methods]. Here we use estimates of the stellar mass from Zhou et al. [31] to split the galaxies in the combined (DR9+DR10) Main sample into four mass/luminosity bins: , each containing 926,232, 1,362,717, 458,886, and 2,152 galaxies, respectively. We note that the estimated masses have not been thoroughly characterized, and thus, they are mostly used to make qualitative statements.
We show the kSZ stacked measurements for each bin in Fig. 6. We detect the individual signals at high significance. The null is , 68.2, 96.2, and 28.3.As can be seen in the figure, we find a strong signature of the evolution of the amplitude with mass: namely, it increases with stellar mass/luminosity, as the optical depth is proportional to the total mass of the halo, i.e., the amplitudes of the stellar-mass bin samples differ roughly by the ratios of their respective mean stellar masses, since (this breaks down as we go to higher halo masses, as the stellar-to-halo mass ratio decreases). As the mean stellar masses in each bin are 1.3, 2.3, 5.0 and 35 , respectively, we see that this statement holds well. We note that the relative differences between the amplitudes are independent of the calibration of the cross-correlation coefficient and the rms of the velocity. We investigate this in detail in simulations of the full DESI data (matching footprint, redshift distribution and halo occupation distribution of the sample) and find that the assumption of scale and mass independence of holds very well: the largest discrepancy we find across all radial bins when testing both the scale and mass independence is 0.2. In addition, as expected, the curves become shallower at small apertures, as the mass increases. This is particularly well seen for the least massive and most massive stellar mass bins, which display the steepest and shallowest profiles, respectively. This is the case because massive halos retain more gas within the virial radius, as they have deeper potential wells compared with their less massive counterparts. At large apertures, the points are extremely strongly correlated, and it is imprudent to draw any definite conclusions on this issue (see App. B). We leave a detailed analysis of the effect of baryons on weak lensing probes, as measured in this paper, to future work.
Appendix F Comparison with CMASS measurement